THE SUBMILLIMETER AND MICROWAVE SPECTRA OF HZnCl (X ^{1}Σ^{+})

\underline{R.L. PULLIAM}, M. SUN, M.A. FLORY, L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectra of HZnCl (^{1}Σ^{+}) has been studied using submillimeter-wave direct-absorption and Fourier transform microwave (FTMW) methods. This species was produced by the reaction of zinc vapor and Cl_{2} in the presence of H_{2} and a d.c. discharge for the millimeter studies. For the FTMW measurements, HZnCl was made by the reaction of chlorine gas and Zn(CH_{3})_{2} in a d.c. discharge created in the supersonic nozzle. The spectra of seven isotopologues were recorded including the deuterium substituted species. We were able to obtain the rotational constants of B_{0} and D_{0} for each isotopologue. Using the FTMW data, we observed the hyperfine components of the lower J transitions and determined the chlorine quadrupole constant. For HZnCl, the average Zn-Cl bond length was determined to be 2.083153(21) Å with the H-Zn bond length equal to 1.51909(38) Å. From HZnCl to its open shelled counterpart ZnCl, the bond length was found to contract by 0.05 Å suggesting that the hydrogen stabilizes the molecule.